Trico Company set the following standard unit costs for its single product.

Direct materials (30 Ibs. @ $4.80 per Ib.) $ 144.00
Direct labor (7 hrs. @ $14 per hr.) 98.00
Factory overhead—variable (7 hrs. @ $6 per hr.) 42.00
Factory overhead—fixed (7 hrs. @ $9 per hr.) 63.00
Total standard cost $ 347.00


The predetermined overhead rate is based on a planned operating volume of 80% of the productive capacity of 67,000 units per quarter. The following flexible budget information is available.

Operating Levels
70% 80% 90%
Production in units 46,900 53,600 60,300
Standard direct labor hours 328,300 375,200 422,100
Budgeted overhead
Fixed factory overhead $ 3,376,800 $ 3,376,800 $ 3,376,800
Variable factory overhead $ 1,969,800 $ 2,251,200 $ 2,532,600


During the current quarter, the company operated at 90% of capacity and produced 60,300 units of product; actual direct labor totaled 399,200 hours. Units produced were assigned the following standard costs.

Direct materials (1,809,000 Ibs. @ $4.80 per Ib.) $ 8,683,200
Direct labor (422,100 hrs. @ $14 per hr.) 5,909,400
Factory overhead (422,100 hrs. @ $15 per hr.) 6,331,500
Total standard cost $ 20,924,100


Actual costs incurred during the current quarter follow.

Direct materials (1,711,000 Ibs. @ $6.30 per lb.) $ 10,779,300
Direct labor (399,200 hrs. @ $11.20 per hr.) 4,471,040
Fixed factory overhead costs 2,813,700
Variable factory overhead costs 3,084,100
Total actual costs $ 21,148,140

Required:
1. Compute the direct materials cost variance, including its price and quantity variances.

2. Compute the direct labor cost variance, including its rate and efficiency variances.

3. Compute the overhead controllable and volume variances.
Answer #1

Solution 1:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
1711000 $6.30 $10,779,300.00 1711000 $4.80 $8,212,800.00 1809000 $4.80 $8,683,200.00
$2,566,500.00 U $470,400.00 F
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $2,566,500.00 U
Direct material quantity variance $470,400.00 F
Direct material cost variance $2,096,100.00 U

Solution 2:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
399200 $11.20 $4,471,040.00 399200 $14.00 $5,588,800.00 422100 $14.00 $5,909,400.00
$1,117,760.00 F $320,600.00 F
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $1,117,760.00 F
Direct Labor Efficiency variance $320,600.00 F
Direct labor cost variance $1,438,360.00 F

Solution 3:

Controllable Variance
Actual overhead $5,897,800.00
Budgeted overhead $5,909,400.00
Controllable variance $11,600.00 F
Fixed overhead volume variance
Budgeted fixed overhead $3,376,800.00
Fixed overhead cost applied $3,798,900.00
Fixed overhead volume variance $422,100.00 F